OpenStudy (erinweeks):

PLEASE SOMEONE HELP ! Allie has a birthday present for her brother. The length of the box that the present comes in is two times longer than the width of the box. The height of the box is 5 inches and its volume is 160 cubic inches. What is the length and width of the box?

OpenStudy (ajprincess):

Let the length of the box be l , width be w and height be h. l=2w. Volume=l*w*h 160=l*w*5 160=2w*w*5 Can u solve for w nw?

OpenStudy (ajprincess):

@ErinWeeks can u do?